# Dispelling the Mirage of Progress in Offline MARL through Standardised Baselines and Evaluation

### v1

@@ 4 Standardising Baselines and Evaluation
While notable progress has been made standardising offline MARL datasets, we maintain that inconsistent use of baselines remains an overlooked issue in the field. Indeed, to highlight this, we conducted extensive benchmarking in Section 3. Now, to address the issue, we release our implementations of BC, IQL+CQL, IDDPG+BC, and MADDPG+CQL, as high-quality baselines for future research. Our baselines come with three main advantages. First, their demonstrated correctness, where we have shown they perform as well as, or better than, most algorithms in the literature on a wide range of datasets. Second, our baselines are easy to parse while also being highly performant on hardware accelerators. The core algorithm logic of our baselines is contained within a single file, which makes it easy for researchers to read, understand, and modify. In addition, all of the training code can be jit (just-in-time) compiled to XLA, making it very fast on hardware accelerators. Furthermore, we use the hardware accelerated replay buffers from Flashbax (Toledo et al., 2023), delivering performance gains by speeding up the time to sample from datasets. The third advantage is their compatibility with OG-MARL datasets, which comes “out of the box”, offering the widest compatibility with offline MARL datasets in the literature. As a foundation for future work, we provide tabular performance values across multiple scenarios for these algorithms, in Table 5. Furthermore, all raw training results can be viewed and downloaded, and are linked to in the appendix.
@@
Recent efforts have been made to address the issue in the online setting (Gorsane et al., 2022; Agarwal et al., 2022). However, similar efforts are still absent in the offline setting, as discussed in Section 2. Following in the spirit of this work, we propose a standardised evaluation methodology for offline MARL, given in the blue box on Page 4, which we believe will significantly improve research outcomes, if adopted.
@@
[hp] A Standardised Evaluation Protocol for Cooperative Offline MARL Choosing the settings to evaluate on: Select at least 2-3 different environments on which to test. For example, evaluating on both SMAC and MAMuJoCo is the most common combination. We encourage additionally evaluating on environments beyond these two, to avoid overfitting to them. Do not use non-standard environment configurations without explicitly stating so. For each environment choose at least 3-4 different scenarios. If the environment creators provide a minimal set of recommended scenarios, use those. Alternatively, focus on scenarios that are common in prior literature. Choose a range of dataset quality types. We recommend at least including a "good" dataset where the majority of samples are from good policies and a "mixed" dataset where samples come from a wide range of policies including good, medium and poor ones. Try to use common existing datasets from the literature (Formanek et al., 2023). If you include your own dataset, provide a clear reason for why, and make it easily accessible to the community. Choosing the baselines to compare to: Choose at least 3-4 relevant baselines to compare to. Usually include behaviour cloning, especially on good datasets where it can be challenging to beat. Try to use common and existing implementations of baselines from the literature. If you include your own novel baseline, make the code available and easy to run for future comparisons. It is not sufficient to only share the code for the novel algorithm being proposed. Choosing the training and evaluation parameters: For each environment, set a training budget and keep it constant across algorithms. For example, we used 100k training updates on SMAC and 200k on MAMuJoCo. If possible, do regular evaluations during training so that you can plot a training curve for analysis at the end of training. Do not use information from these evaluations to influence a training run online, as this would violate the assumptions around the training being offline. We recommend at every evaluation step unrolling for 32 episodes and reporting the average episode return. As per Agarwal et al. (2022), you should repeat each run across 10 random seeds. Reporting your results: Report the final evaluation result, averaged across all 10 seeds, along with the standard deviation. If doing regular evaluations during training, also report the average and maximum episode return during training as per Papoudakis et al. (2021) and plot sample efficiency curves as per Gorsane et al. (2022). Use appropriate statistical significance calculations to report on whether your algorithms significantly outperform the baselines (Papoudakis et al., 2021; Agarwal et al., 2022). In addition to reporting results on a per-dataset basis, also report aggregated results across scenarios and dataset types. Results can be aggregated by first normalising them as per Fu et al. (2021) and then applying the utilities from MARL-eval (Gorsane et al., 2022) (e.g. performance profile plots, see Figure 4).
@@ Appendix D Converted Datasets
Table 15: Links to Vault datasets. Those from Formanek et al. [2023] were already in the Vault format; the others were converted.
Paper Environment Scenario Link Formanek et al. [2023] SMACv1 3m https://s3.kao.instadeep.io/offline-marl-dataset/vaults/3m.zip 8m https://s3.kao.instadeep.io/offline-marl-dataset/vaults/8m.zip 5m_vs_6m https://s3.kao.instadeep.io/offline-marl-dataset/vaults/5m_vs_6m.zip 2s3z https://s3.kao.instadeep.io/offline-marl-dataset/vaults/2s3z.zip 3s5z_vs_3s6z https://s3.kao.instadeep.io/offline-marl-dataset/vaults/3s5z_vs_3s6z.zip 2c_vs_64zg https://s3.kao.instadeep.io/offline-marl-dataset/vaults/2c_vs_64zg.zip MAMuJoCo 2ant https://s3.kao.instadeep.io/offline-marl-dataset/vaults/2ant.zip 2halfcheetah https://s3.kao.instadeep.io/offline-marl-dataset/vaults/2halfcheetah.zip 4ant https://s3.kao.instadeep.io/offline-marl-dataset/vaults/4ant.zip Pan et al. [2022] MAMuJoCo 2halfcheetah https://s3.kao.instadeep.io/offline-marl-dataset/omar/2halfcheetah.zip MPE simple-spread https://s3.kao.instadeep.io/offline-marl-dataset/omar/simple_spread.zip Shao et al. [2023] SMACv1 2s3z https://s3.kao.instadeep.io/offline-marl-dataset/cfcql/2s3z.zip 3s_vs_5z https://s3.kao.instadeep.io/offline-marl-dataset/cfcql/3s_vs_5z.zip 5m_vs_6m https://s3.kao.instadeep.io/offline-marl-dataset/cfcql/5m_vs_6m.zip 6h_vs_8z https://s3.kao.instadeep.io/offline-marl-dataset/cfcql/6h_vs_8z.zip Wang et al. [2023] SMACv1 corridor https://s3.kao.instadeep.io/offline-marl-dataset/omiga/corridor.zip 2c_vs_64zg https://s3.kao.instadeep.io/offline-marl-dataset/omiga/2c_vs_64zg.zip 5m_vs_6m https://s3.kao.instadeep.io/offline-marl-dataset/omiga/5m_vs_6m.zip 6h_vs_8z https://s3.kao.instadeep.io/offline-marl-dataset/omiga/6h_vs_8z.zip MAMuJoCo 2ant https://s3.kao.instadeep.io/offline-marl-dataset/omiga/2ant.zip 3hopper https://s3.kao.instadeep.io/offline-marl-dataset/omiga/3hopper.zip 6halfcheetah https://s3.kao.instadeep.io/offline-marl-dataset/omiga/6halfcheetah.zip
@@
To download and view the profile of a dataset, you can easily use the utilities in the attached code. For example,
1import jax.numpy as jnp
2from jax.tree_util import tree_map
3from utils.offline_dataset import (
4 download_and_unzip_vault,
5 analyse_vault,
6)
7
8scenario = "5m_vs_6m"
9sources = ["smac_v1", "smac_v1_omiga", "smac_v1_cfcql"]
10
11for source in sources:
12 download_and_unzip_vault(source, scenario)
13 returns_dict = analyse_vault(
14 f"{source}/{scenario}.vlt", visualise=True,
15 )
16 print(f"Source: {source}")
17 print(tree_map(lambda x: [
18 float(jnp.mean(x)), float(jnp.std(x))
19 ], returns_dict))
which yields the violin plots in Figure 9, and the summary statistics in Table 16.
(a) Formanek et al. [2023]
(b) Wang et al. [2023]
(c) Shao et al. [2023]
Figure 9: Statistical profiles of datasets from the literature, on the SMACv1 5m_vs_6m scenario.
Table 16: Summary statistics of datasets from the literature, on the SMACv1 5m_vs_6m scenario.
Paper Environment Scenario Quality Mean ± Std Formanek et al. [2023] SMACv1 5m_vs_6m Good $16.6\pm 4.7$ Medium $12.6\pm 5.1$ Poor $7.5\pm 1.5$ Wang et al. [2023] SMACv1 5m_vs_6m Good $20.0\pm 0.0$ Medium $11.0\pm 0.6$ Poor $8.5\pm 1.2$ Shao et al. [2023] SMACv1 5m_vs_6m Expert $18.2\pm 3.8$ Medium $12.0\pm 4.4$ Medium-Replay $9.0\pm 2.6$ Mixed $15.1\pm 5.1$

### v2

@@ 4 Standardising Baselines and Evaluation
While notable progress has been made standardising offline MARL datasets, we maintain that inconsistent use of baselines remains an overlooked issue in the field. Indeed, to highlight this, we conducted extensive benchmarking in Section 3. Now, to address the issue, we release our implementations of BC, IQL+CQL, IDDPG+BC, and MADDPG+CQL, as high-quality baselines for future research. Our baselines come with three main advantages. First, their demonstrated correctness, where we have shown they perform as well as, or better than, most algorithms in the literature on a wide range of datasets. Second, our baselines are easy to parse while also being highly performant on hardware accelerators. The core algorithm logic of our baselines is contained within a single file, which makes it easy for researchers to read, understand, and modify. In addition, all of the training code can be jit (just-in-time) compiled to XLA, making it very fast on hardware accelerators. Furthermore, we use the hardware accelerated replay buffers from Flashbax (Toledo et al., 2023), delivering performance gains by speeding up the time to sample from datasets. The third advantage is their compatibility with OG-MARL datasets, which comes “out of the box”, offering the widest compatibility with offline MARL datasets in the literature  [^2] Datasets and baselines can be accessed at https://github.com/instadeepai/og-marl. As a foundation for future work, we provide tabular performance values across multiple scenarios for these algorithms, in Table 5. Furthermore, all raw training results can be viewed and downloaded, and are linked to in the appendix.
@@
Recent efforts have been made to address the issue in the online setting (Gorsane et al., 2022; Agarwal et al., 2022). However, similar efforts are still absent in the offline setting, as discussed in Section 2. Following in the spirit of this work, we propose a set of evaluation guidelines for offline MARL, given in the blue box on Page 4, which we believe will significantly improve research outcomes, if adopted.
@@
[hp] Evaluation Guidelines for Cooperative Offline MARL Choosing the settings to evaluate on: Select at least 2-3 different environments on which to test. For example, evaluating on both SMAC and MAMuJoCo is the most common combination. We encourage additionally evaluating on environments beyond these two, to avoid overfitting to them. Do not use non-standard environment configurations without explicitly stating so. For each environment choose at least 3-4 different scenarios. If the environment creators provide a minimal set of recommended scenarios, use those. Alternatively, focus on scenarios that are common in prior literature. Choose a range of dataset quality types. We recommend at least including a "good" dataset where the majority of samples are from good policies and a "mixed" dataset where samples come from a wide range of policies including good, medium and poor ones. Try to use common existing datasets from the literature (Formanek et al., 2023). If you include your own dataset, provide a clear reason for why, and make it easily accessible to the community. Choosing the baselines to compare to: Choose at least 3-4 relevant baselines to compare to. Usually include behaviour cloning, especially on good datasets where it can be challenging to beat. Try to use common and existing implementations of baselines from the literature. If you include your own novel baseline, make the code available and easy to run for future comparisons. It is not sufficient to only share the code for the novel algorithm being proposed. Choosing the training and evaluation parameters: For each environment, set a training budget and keep it constant across algorithms. For example, we used 100k training updates on SMAC and 200k on MAMuJoCo. If possible, do regular evaluations during training so that you can plot a training curve for analysis at the end of training. Do not use information from these evaluations to influence a training run online, as this would violate the assumptions around the training being offline. We recommend at every evaluation step unrolling for 32 episodes and reporting the average episode return. As per Agarwal et al. (2022), you should repeat each run across 10 random seeds. Reporting your results: Report the final evaluation result, averaged across all 10 seeds, along with the standard deviation. If doing regular evaluations during training, also report the average and maximum episode return during training as per Papoudakis et al. (2021) and plot sample efficiency curves as per Gorsane et al. (2022). Use appropriate statistical significance calculations to report on whether your algorithms significantly outperform the baselines (Papoudakis et al., 2021; Agarwal et al., 2022). In addition to reporting results on a per-dataset basis, also report aggregated results across scenarios and dataset types. Results can be aggregated by first normalising them as per Fu et al. (2021) and then applying the utilities from MARL-eval (Gorsane et al., 2022) (e.g. performance profile plots, see Figure 4).
@@ Appendix D Converted Datasets
Table 15: Links to Vault datasets converted from the literature.
Paper Environment Scenario Link Pan et al. [2022] MAMuJoCo 2halfcheetah https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omar/mamujoco/2halfcheetah.zip MPE simple-spread https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omar/mpe/simple_spread.zip simple-tag https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omar/mpe/simple_tag.zip simple-world https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omar/mpe/simple_world.zip Shao et al. [2023] SMACv1 2s3z https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/cfcql/smac_v1/2s3z.zip 3s_vs_5z https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/cfcql/smac_v1/3s_vs_5z.zip 5m_vs_6m https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/cfcql/smac_v1/5m_vs_6m.zip 6h_vs_8z https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/cfcql/smac_v1/6h_vs_8z.zip Wang et al. [2023] SMACv1 corridor https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omiga/smac_v1/corridor.zip 2c_vs_64zg https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omiga/smac_v1/2c_vs_64zg.zip 5m_vs_6m https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omiga/smac_v1/5m_vs_6m.zip 6h_vs_8z https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omiga/smac_v1/6h_vs_8z.zip MAMuJoCo 2ant https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omiga/mamujoco/2ant.zip 3hopper https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omiga/mamujoco/3hopper.zip 6halfcheetah https://huggingface.co/datasets/InstaDeepAI/og-marl/resolve/main/prior_work/omiga/mamujoco/6halfcheetah.zip
@@
The statistical profiles (histograms and violin plots) of the converted datasets are available in the respective folders on HuggingFace: https://huggingface.co/datasets/InstaDeepAI/og-marl/tree/main/prior_work. Additionally, summary statistic values can easily be accessed using the demonstrative notebook at: https://github.com/instadeepai/og-marl/blob/main/examples/dataset_analysis_demo.ipynb.